OpenStudy (anonymous):

help me solve this question. The tens' digit of a two-digit number is 3 more than the units' digit. the number is 8 more than six times the sum of the digits. find the number

OpenStudy (anonymous):

please help me set up the equation

OpenStudy (anonymous):

let's take y as the tens and x as the units. We have : y = x+3 and y*10 + x = 6 * (x+y) +8

OpenStudy (anonymous):

can you manage from here?